# MVGT: A Multi-view Graph Transformer Based on Spatial Relations for EEG Emotion Recognition

### v1

@@ I Introduction
The spatial structure of the brain also contains rich emotional information. Emotional states may involve distributed circuits rather than considering a single brain region in isolation[8]. Asymmetry between the left and right hemispheres can reflect changes in valence and arousal[9]. Recent studies have highlighted the importance of utilizing spatial domain information. Li et al.[10] introduced recurrent neural networks to learn the asymmetric differences between the left and right hemispheres. Li et al.[11] also utilized hierarchical neural networks to learn both regional and global information of spatial-temporal EEG features. Graph Neural Networks (GNN) are emerging as a powerful tool for analyzing spatial information in EEG emotion recognition. Song et al.[12] dynamically learned relationships between EEG channels using graph convolutional networks (GCN). Zhong et al.[13] incorporated asymmetry of the left and right hemispheres into the adjacency matrix to model graph structure. Li et al.[14] also utilized adaptive graph convolutional networks that integrate multi-domain information to learn relationships between channels. Ding et al.[15] incorporated lobe information as prior knowledge into the GNN. Jiang et al.[16] proposed an elastic graph Transformer to extract emotional information. Although these methods have achieved excellent performance in emotion recognition tasks, they have a common issue: they all rely on GNNs based on neighborhood aggregation schemes which may pose potential risks such as over-smoothing[17, 18, 19], under-reaching[20], and over-squashing[21]. Additionally, these methods do not comprehensively consider the geometric and anatomical structure information of the brain.
The main contributions of this paper are as follows:
@@ II-A EEG-based emotion recognition
Along with the development of deep learning, increasingly complex models with rich expressive abilities have emerged and have been extensively utilized in EEG signal analysis. Zheng et al.[5] employed deep belief networks to analyze important frequency domain components and effective channels based on the learned parameters. Song et al.[12] used a graph convolutional method based on Chebyshev polynomials[24] to dynamically learn the representations of EEG signals. Zhong et al.[13] innovatively incorporated the asymmetric information of the hemispheres as prior knowledge into the adjacency matrix in 3D space and used GCN to dynamically learn the inter-channel correlations. The reasonable combination of the multi-domain information contributes to improving the accuracy in the emotion recognition task. Li et al.[14] proposed an adaptive graph convolutional network that integrates the temporal domain, frequency domain, and functional connectivity. Ding et al.[15], inspired by neuroscience research, combined intra-region convolution and inter-region convolution based on brain lobe regions to learn brain cognitive patterns. Jiang et al.[16] utilized the advantages of GCN in the spatial domain and Transformer in the temporal domain to improve the accuracy of emotion classification.
@@ II-B Graph Transformer
The GNNs used in the above methods are based on neighborhood aggregation schemes. However, classical GNNs based on message passing (MPGNNs) may lead to over-smoothing[17, 18, 19], under-reaching[20], and may also fail to fit long-range signals due to over-squashing[21], which limit the expressive power of the model. Graph transformers (GTs) alleviate such effects as they have a global receptive field[25]. However, without sufficiently expressive structural and positional encodings, GTs cannot capture effective graph structures[26]. Dwivedi et al.[27] utilized eigenvectors of graph Laplacian as position encodings in fully connected Graph Transformers and integrated edge features into the attention mechanism. Building on this, SAN[28] used a full Laplacian spectrum to learn the positional encodings for each node. Graphormer[29, 30] employed node centrality and node distance metric to implement structural and relative positional encodings, achieving state-of-the-art performance on molecular prediction datasets. In EEG emotion recognition, Li et al.[31] innovatively combined a masked autoencoder based on self-supervised learning with a CNN-Transformer hybrid structure, effectively improving classification accuracy. However, this method only used sine-cosine positional encodings, limiting the Transformer’s ability to learn spatial information.
@@ III-A Graph Neural Network (GNN)
Let $G=\left(V,E\right)$ define a graph, where $V=\left\{v_{1},v_{2},\cdots,v_{n}\right\}$ represents the nodes in the graph, and $E=\left\{e_{1},e_{2},\cdots,e_{m}\right\}$ is the edges between the nodes. The representation of node $v_{i}$ is denoted as $x_{i}\in\mathbb{R}^{d}$. Most existing GNNs[32, 17, 33, 34, 35] adopt neighborhood aggregation schemes, iteratively aggregating representations of its first or higher-order neighbors, followed by using backpropagation (BP) to learn task-driven feature encodings. We define the representation of node $v_{i}$ at the $l$ iteration as $h_{i}^{(l)}$ and define $h_{i}^{(0)}=x_{i}$. The $l$-th iteration can be represented as:
$a_{i}^{(l)}$ $=\text{AGGREGATE}^{(l)}\left(\left\{\varphi_{\theta}(h_{j}^{(l)},e_{ji}^{(l)}):j\in\mathcal{N}(v_{i})\right\}\right)$ (1)
$h_{i}^{(l)}$ $=\text{UPDATE}^{(l)}\left(h_{i}^{(l-1)},a_{i}^{(l)}\right)$ (2)
where $\varphi_{\theta}$ represents a differentiable function used for feature transformation of node and edge information. The $\mathcal{N}(v_{i})$ is the set of neighbors of $v_{i}$. The AGGREGATE function is used to aggregate the transformed representation using a differentiable, permutation invariant function, (such as mean, sum, max, etc.). The goal of UPDATE function is to integrate the information from neighbors into the node representation. For graph classification, the READOUT operation is typically used to obtain a representation of the entire graph, which is then fed into a classifier to determine the graph label.
@@ III-B Graphormer
The Transformer[36] is undeniably one of the most popular deep neural network architectures today, driving significant advancements in natural language processing and computer vision. With its global receptive field and multi-head attention mechanism, Transformer can extract global semantic correlations between tokens in multiple feature subspaces, effectively enhancing the model’s expressive power. From the perspective of GNNs, Transformer can be interpreted as a GNN acting on a fully connected graph. Therefore, it is reasonable and feasible to use Transformer to address tasks on graph data. The ability to properly incorporate the structural information of graphs into the model is the key for leveraging its expressive power. Graphormer[29, 30] can go beyond classical MPGNNs in expressive power and achieves state-of-the-art performance on large molecular benchmarks. Graphormer incorporates centrality encoding into the graph data and integrates spatial encoding, edge encoding into the attention mechanism, which can be expressed as:
@@
where the bias term $b_{\phi(v_{i},v_{j})}$ can adaptively adjust the correlations between $v_{i}$ and $v_{j}$. The $c_{ij}$ represents the edge encoding on the shortest path.
@@ IV-A Problem Definition
EEG signals can be represented as a two-dimensional matrix with respect to channels and time. Given that channels exhibit spatial structure, they can be structured into fully connected graph data $G=\left(V,E\right)$, where $V$ denotes the nodes in the graph, representing EEG channels, and $E$ denotes the edges, representing the connections between channels. The features of the nodes are denoted by $X=\left(x_{1},x_{2},\cdots,x_{n}\right)\in\mathbb{R}^{n\times d}$, where $n=|V|$ represents the number of nodes and $d$ represents the feature dimension.
@@ IV-B Temporal Embedding
EEG signals have high temporal resolution and contain rich temporal information. Because of the multi-electrode acquisition method, EEG signals can be regarded as multivariate time series. When processing time series, the embedding of temporal information are crucial. EmoGT[16] treats the features of different channels at the same time points as tokens and employs an attention mechanism to extract temporal correlations between them. Due to the different anisotropic volume conduction characteristics[37] in human brain tissues, there may be temporal delays between different channels, which in turn leads to time-unaligned events at a single moment thus causing performance degradation. MD-AGCN[14] utilizes convolutional neural networks to extract temporal information along the time axis from continuous EEG segments, with the receptive field limited by the size of the convolution kernel. Inspired by iTransformer[38], we broaden the receptive field by considering the entire time series as an embedded token rather than a single time point. First, following the methods of MD-AGCN and EmoGT, we use overlapping sliding windows of size $T$ to segment EEG signals along the time axis and use these segments as tokens, which are then fed into the attention module in the form of continuous segments. After processing with sliding windows, we obtain $\bm{X}\in\mathbb{R}^{S\times n\times Tf}$, where $S$ denotes the number of continuous EEG segments, $n$ is the number of channels, and $f$ denotes the dimension of frequency domain features.
@@ IV-C1 Brain Region Encoding
Neuroscience research demonstrated that the activation of a specific brain region often leads to the concurrent activation of related brain regions responsible for the same high-level cognition[40]. In EEG emotion recognition, incorporating relevant neuroscience findings can typically enhance recognition accuracy. RGNN[13] integrates the asymmetry of neural activity between the left and right hemispheres as prior knowledge into the adjacency matrix, effectively enhancing recognition accuracy. BiHDM[10] improves emotion pattern recognition performance by learning the differences between the left and right hemispheres. LGGNet[15] divides EEG channels into different regions and combines local intra-region convolution with global inter-region convolution, achieving good results on the DEAP[41] dataset. With reference to the three divisions of LGGNet, we adopt four brain region divisions, which divide the EEG channels into different regions based on a prior knowledge, aiming to incorporate the brain region information into the model.
We divide the regions based on the anatomical structure of the brain and implement LOBE scheme.
To further investigate the expressive power of brain region encoding, we conduct a detailed division of brain lobes according to the 10-20 system based on electrode positions, employing the GENERAL scheme.
Asymmetric EEG activity in the frontal lobe can be utilized for discriminating valence changes[9]. The left frontal lobe exhibits a stronger correlation with joy and happy, while the right frontal lobe is more strongly correlated with fear and sadness. Thus we further divide the frontal lobe region into two symmetrical regions to obtain the FRONTAL scheme.
According to the symmetry of brain structure[42], we make a finer division of the brain lobe regions, defining the HEMISPHERE scheme.
The four modes mentioned above are showed in Fig. 2. In terms of specific implementation, we assign each electrode a brain region tag, then project the tags into an embedding space using a learnable projection function, and simply add the embeddings to the node features. The encoding of node $i$ is represented as follows:
@@ IV-C2 Geometric Structure Encoding
In the real world, the human reasoning process considers not only the semantic relationships between objects but also their spatial relations. EEG channels have a 3D structure, and the functional connectivity between these channels lack precise definitions. Therefore, we represent the relationships between EEG channels as a fully connected directed graph structure. The Euclidean distance between channels is calculated using their coordinates to learn the spatial correlations between nodes. Firstly, let $\phi(i,j)$ represents the Euclidean distance between node $i$ and node $j$, and encode $\phi(i,j)$ using a set of Gaussian basis functions [43, 30]. Let $\bm{b}_{k}\in\mathbb{R}^{n\times n}$ denotes one of the Gaussian basis functions. The element $(i,j)$ of this function can be expressed as:
@@
$\bm{h}_{i}^{(0)}=\bm{x}_{i}\bm{W}_{\mathcal{X}}+\bm{z}_{i}\bm{W}_{\mathcal{Z}}+\bm{r}_{i},\ \bm{z}_{i}=\sum_{j=1}^{n}\bm{B}_{i,j,k},$ (7)
where $i$ denotes the node index, and $\bm{W}_{\mathcal{Z}}\in\mathbb{R}^{K\times d}$ is a learnable projection function. Additionally, we incorporate the spatial encoding as a bias term into the softmax attention, which will help the model properly capture the spatial correlations.
Our proposed spatial encoding matrix is directed, which is inconsistent with the assumption of a symmetric adjacency matrix[16, 13]. Using directed connections provides the model with greater expressive power because the correlation between node pairs $\left(i,j\right)$ and $\left(j,i\right)$ may differ. Since we assume nodes are fully connected, we avoid specific assumptions about inter-channel correlations and learn the functional correlations between nodes through encoding. Let $l$ denote the model depth, and $i$ denote the index of multi-head attention. Therefore, the brain functional encoding can be represented as:
$\bm{A}^{l,i}=\text{Softmax}\left(\frac{\left(\bm{H}^{l}\bm{W}_{\mathcal{Q}}^{l,i}\right)\left(\bm{H}^{l}\bm{W}_{\mathcal{K}}^{l,i}\right)^{T}}{\sqrt{d^{l}}}+\bm{B}\bm{W}_{\mathcal{B}}^{i}\right),$ (8)
where $\bm{W}_{\mathcal{Q}}^{l,i}$, $\bm{W}_{\mathcal{K}}^{l,i}$ and $\bm{W}_{\mathcal{B}}^{i}$ are learnable parameters, and $d^{l}$ denotes the feature dimension size of the $l$-th layer. This encoding method integrates temporal, frequency, and spatial domain features into the model, enhancing its expressive power. We compute the attention scores between nodes using embedded vectors, representing the semantic correlations between different nodes from multiple perspectives. Finally, the attention scores are added to the spatial geometric encoding to obtain the correlations between channels.

### v2

@@ I Introduction
The spatial structure of the brain also contains rich emotional information. Emotional states may involve distributed circuits rather than considering a single brain region in isolation [8]. Asymmetry between the left and right hemispheres can reflect changes in valence and arousal [9]. Recent studies have highlighted the importance of utilizing spatial domain information. Li et al. [10] introduced recurrent neural networks to learn the asymmetric differences between the left and right hemispheres. Li et al. [11] also utilized hierarchical neural networks to learn both regional and global information of spatial-temporal EEG features. Graph Neural Networks (GNN) are emerging as a powerful tool for analyzing spatial information in EEG emotion recognition. Song et al. [12] dynamically learned relationships between EEG channels using graph convolutional networks (GCN). Zhong et al. [13] incorporated asymmetry of the hemispheres into the adjacency matrix to model graph structure. Li et al. [14] also utilized adaptive GCNs that integrate multi-domain information to learn relationships between channels. Ding et al. [15] incorporated lobe information as prior knowledge into the GNN. Jiang et al. [16] proposed an elastic graph Transformer to extract emotional information. Although these methods have achieved excellent performance in emotion recognition, they have a common issue: they all rely on GNNs based on neighborhood aggregation schemes which may pose potential risks such as over-smoothing [17, 18, 19], under-reaching [20], and over-squashing [21]. Additionally, these methods do not comprehensively consider the geometric and anatomical structure information of the brain.
The main contributions of this paper lie in three aspects:
@@ II-A EEG-based Emotion Recognition
Along with the development of deep learning, increasingly complex models with rich expressive abilities have emerged and have been extensively utilized in EEG signal analysis. Zheng et al. [5] employed deep belief networks to analyze important frequency domain components and effective channels based on the learned parameters. Song et al. [12] used a GCN method based on Chebyshev polynomials [24] to dynamically learn the representations of EEG signals. Zhong et al. [13] innovatively incorporated the asymmetric information of the hemispheres as prior knowledge into the adjacency matrix in 3D space. The reasonable combination of the multi-domain information contributes to improving the accuracy in the emotion recognition task. Li et al. [14] proposed an adaptive GCN that integrates the temporal domain, frequency domain, and functional connectivity. Ding et al. [15], inspired by neuroscience research, combined intra-region convolution and inter-region convolution based on brain lobe regions to learn cognitive patterns. Jiang et al. [16] utilized the advantages of GCN in the spatial domain and Transformer in the temporal domain to improve the accuracy of emotion classification.
@@ II-B Graph Transformer
The GNNs used in the above methods are based on neighborhood aggregation schemes. However, classical GNNs based on message passing (MPGNNs) may lead to over-smoothing [17, 18, 19], under-reaching [20], and may also fail to fit long-range signals due to over-squashing [21], which limit the expressive power of the model. Graph transformers (GTs) alleviate such effects as they have a global receptive field [25]. Nevertheless, without sufficiently expressive structural and positional encodings, GTs cannot capture effective graph structures [26]. Dwivedi et al. [27] utilized eigenvectors of graph Laplacian as position encodings in fully connected Graph Transformers and integrated edge features into the attention mechanism. Building on this, SAN [28] used a full Laplacian spectrum to learn the positional encodings for each node. Graphormer [29, 30] employed node centrality and node distance metric to implement structural and relative positional encodings, achieving state-of-the-art performance on challenging graph datasets. In EEG emotion recognition, Li et al. [31] innovatively combined a masked autoencoder based on self-supervised learning with a CNN-Transformer hybrid structure, effectively improving classification accuracy. However, this method only used the sine-cosine positional encoding, limiting the Transformer’s ability to learn spatial information.
@@ III-A Graph Neural Network (GNN)
Let $G=\left(V,E\right)$ define a graph, where $V=\left\{v_{1},v_{2},\cdots,v_{n}\right\}$ represents the nodes in the graph, and $E=\left\{e_{1},e_{2},\cdots,e_{m}\right\}$ is the edges between the nodes. The representation of node $v_{i}$ is denoted as $x_{i}\in\mathbb{R}^{d}$. Most existing GNNs [32, 17, 33, 34, 35] adopt neighborhood aggregation schemes, iteratively aggregating representations of its first or higher-order neighbors, followed by using backpropagation (BP) to learn task-driven feature encodings. We define the representation of node $v_{i}$ at the $l$ iteration as $h_{i}^{l}$ and define $h_{i}^{0}=x_{i}$. The $l$-th iteration can be represented as:
$a_{i}^{l}$ $=\text{AGGREGATE}^{l}\left(\left\{\varphi_{\theta}(h_{j}^{l},e_{ji}^{l}):j\in\mathcal{N}(v_{i})\right\}\right)$ (1)
$h_{i}^{l}$ $=\text{UPDATE}^{l}\left(h_{i}^{l-1},a_{i}^{l}\right)$ (2)
where $\varphi_{\theta}$ represents a differentiable function used for feature transformation of node and edge information. $\mathcal{N}(v_{i})$ is the set of neighbors of $v_{i}$. The AGGREGATE function is used to aggregate the transformed representation using a differentiable, permutation invariant function, (such as mean, sum, max, etc.). The goal of UPDATE function is to integrate the information from neighbors into the node representation. For graph classification, the READOUT operation is typically used to obtain a representation of the entire graph, which is then fed into a classifier to determine the graph label.
@@ III-B Graphormer
The Transformer [36] is undeniably one of the most popular deep neural network architectures today, driving significant advancements in natural language processing and computer vision. From the perspective of GNNs, Transformer can be interpreted as a GNN acting on a fully connected graph. Therefore, it is reasonable and feasible to use Transformer to address tasks on graph data. The ability to properly incorporate the structural information of graphs into the model is the key for leveraging its expressive power. Graphormer [29, 30] can go beyond classical MPGNNs in expressive power and achieves state-of-the-art performance on large molecular benchmarks. Graphormer incorporates centrality encoding into the graph data and integrates spatial encoding, edge encoding into the attention mechanism, which can be expressed as:
@@
where $c_{ij}$ represents the edge encoding on the shortest path and the bias term $b_{\phi(v_{i},v_{j})}$ can adaptively adjust the correlations between $v_{i}$ and $v_{j}$.
@@ IV-A Problem Definition
EEG signals can be represented as a two-dimensional matrix with respect to channels and time. Given that channels exhibit spatial structure, they can be structured into fully connected graph data $G=\left(V,E\right)$, where $V$ denotes the nodes in the graph, representing EEG channels, and $E$ denotes the edges, representing the connections between channels. The features of the nodes at time $t$ are denoted by $\bm{X}_{t}=\left(\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}\right)\in\mathbb{R}^{n\times d}$, where $n=|V|$ represents the number of nodes and $d$ represents the feature dimension.
@@ IV-B Temporal Embedding
EEG signals have high temporal resolution and contain rich temporal information. Because of the multi-electrode acquisition method, EEG signals can be regarded as multivariate time series. When processing time series, the embedding of temporal information are crucial. EmoGT [16] treats the features of different channels at the same time points as tokens and employs an attention mechanism to extract temporal correlations between them. Due to the different anisotropic volume conduction characteristics [37] in human brain tissues, there may be temporal delays between different channels, which in turn leads to time-unaligned events at a single moment thus causing performance degradation. MD-AGCN [14] utilizes convolutional neural networks to extract temporal information along the time axis from continuous EEG segments, with the receptive field limited by the size of the convolution kernel. Inspired by iTransformer [38], we broaden the receptive field by considering the entire time series as an embedded token rather than a single time point. First, following the methods of MD-AGCN and EmoGT, we use overlapping sliding windows of size $T$ to segment EEG signals along the time axis and use these segments as tokens, which are then fed into the attention module in the form of continuous segments. After processing with sliding windows, we obtain $\tilde{\bm{X}}=\left(\tilde{\bm{X}}_{1},\tilde{\bm{X}}_{2},\cdots,\tilde{\bm{X}}_{S}\right)\in\mathbb{R}^{S\times n\times Tf}$, where $S$ denotes the number of continuous EEG segments, $n$ is the number of channels, and $f$ denotes the dimension of frequency domain features.
@@ IV-C1 Brain Region Encoding
Neuroscience research demonstrated that the activation of a specific brain region often leads to the concurrent activation of related brain regions responsible for the same high-level cognition [40]. In EEG emotion recognition, incorporating relevant neuroscience findings can typically enhance recognition accuracy. RGNN [13] integrates the asymmetry of neural activity between the left and right hemispheres as prior knowledge into the adjacency matrix, effectively enhancing recognition accuracy. BiHDM [10] improves emotion pattern recognition performance by learning the differences between the left and right hemispheres. LGGNet [15] divides EEG channels into different regions and combines local intra-region convolution with global inter-region convolution, achieving good results on the DEAP [41] dataset. With reference to the three divisions of LGGNet, we adopt four brain region divisions, which divide the EEG channels into different regions based on a prior knowledge, aiming to incorporate the brain region information into the model. The division schemes consist of: (1) We divide the regions based on the anatomical structure of the brain and implement LOBE scheme. (2) To further investigate the expressive power of brain region encoding, we conduct a detailed division of brain lobes according to the 10-20 system based on electrode positions, employing the GENERAL scheme. (3) Asymmetric EEG activity in the frontal lobe can be utilized for discriminating valence changes [9]. The left frontal lobe exhibits a stronger correlation with joy and happy, while the right frontal lobe is more strongly correlated with fear and sadness. Thus we further divide the frontal lobe region into two symmetrical regions to obtain the FRONTAL scheme. (4) According to the symmetry of brain structure [42], we make a finer division of the brain lobe regions, defining the HEMISPHERE scheme.
The four schemes mentioned above are showed in Fig. 2. In terms of specific implementation, we assign each electrode a brain region tag, then project the tags into an embedding space using a learnable projection function, and simply add the embeddings to the node features. The encoding of node $i$ is represented as follows:
@@ IV-C2 Geometric Structure Encoding
In the real world, the human reasoning process considers not only the semantic relationships between objects but also their spatial relations. EEG channels have a 3D structure, and the functional connectivity between these channels lack precise definitions. Therefore, we represent the relationships between EEG channels as a fully connected directed graph structure. The Euclidean distance between channels is calculated using their coordinates to learn the spatial correlations between nodes. Firstly, let $\phi(i,j)$ represent the Euclidean distance between node $i$ and node $j$, and encode $\phi(i,j)$ using a set of Gaussian basis functions [43, 30]. Let $\bm{b}_{k}\in\mathbb{R}^{n\times n}$ denote one of the Gaussian basis functions. The element $(i,j)$ of this function can be expressed as:
@@
$\bm{h}_{i}^{0}$ $=\bm{x}_{i}\bm{W}_{\mathcal{X}}+\bm{z}_{i}\bm{W}_{\mathcal{Z}}+\bm{r}_{i},\ \bm{z}_{i}=\sum_{j=1}^{n}\bm{B}_{i,j,k},$ (7)
$\bm{B^{\prime}}$ $=\text{Projection}(\bm{B}),$ (8)
where $i$ is the node index, $k$ is the index of basis function and $\bm{W}_{\mathcal{Z}}\in\mathbb{R}^{K\times d}$ is a learnable projection function. $\text{Projection}:\mathbb{R}^{n\times n\times K}\mapsto\mathbb{R}^{n\times n\times M}$ is a nonlinear transformation, where $M$ is the number of attention heads. We incorporate this encoding as a bias term into the softmax attention.
Our proposed spatial encoding matrix is directed, which is inconsistent with the assumption of a symmetric adjacency matrix [16, 13]. Using directed connections provides the model with greater expressive power because the correlation between node pairs $\left(i,j\right)$ and $\left(j,i\right)$ may differ. Since we assume nodes are fully connected, we avoid specific assumptions about inter-channel correlations and learn the functional correlations between nodes through encoding. Let $l$ denote the model depth, and $i$ denote the index of multi-head attention. Therefore, the brain functional encoding of the $s$-th sample can be represented as:
$\bm{A}^{l,i}=\text{Softmax}\left(\frac{\left(\bm{H}^{l}_{s}\bm{W}_{\mathcal{Q}}^{l,i}\right)\left(\bm{H}^{l}_{s}\bm{W}_{\mathcal{K}}^{l,i}\right)^{T}}{\sqrt{d_{h}^{l,i}}}+\bm{B^{\prime}}^{i}\right),$ (9)
where the projections are learnable parameters $\bm{W}_{\mathcal{Q}}^{l,i}\in\mathbb{R}^{d\times d_{h}}$ and $\bm{W}_{\mathcal{K}}^{l,i}\in\mathbb{R}^{d\times d_{h}}$, and $d_{h}^{l,i}$ is the second dimension of $\bm{W}_{\mathcal{K}}^{l,i}$. The scalar $d_{h}$ represents the number of the hidden dimension. This encoding method integrates temporal, frequency, and spatial domain features into the model, enhancing its expressive power. We compute the attention scores between nodes using embedded vectors, representing the semantic correlations between different nodes from multiple perspectives. Finally, the attention scores are added to the spatial geometric encoding to obtain the correlations between channels.